OpenStudy (anonymous):

log base5 20 * log base20 125. How do I find the exact value of this expression?

OpenStudy (phi):

log_b(x) = log_a(x)/log_a(b)

OpenStudy (phi):

rewrite log_20(125) in terms of log_5 using the above

OpenStudy (campbell_st):

use change of base \[\log_{a}(b) = \frac{\ln(b)}{\ln(a)}\] it doesn't matter if you use base 10 or base e logs so your problems becomes \[\frac{\ln(20}{\ln(5)} \times \frac{\ln(125)}{\ln(20)}\] now you can use some log laws and find common factors \[\frac{\ln(20)}{\ln(5)} \times \frac{\ln(5^3)}{\ln(20)}\] good luck with completing the answer.